=================================================================== Tohban report: 2009-03-22 checked data from: 2009-03-15/00:00 UT to 2009-03-20/24:00 UT =================================================================== Data quality/collection status ============= FGM: A: 3/15-3/20: full data coverage B: 3/15-3/20: full data coverage C: 3/15-3/20: full data coverage D: 3/15-3/20: data gaps 3/17 12:55 - 3/17 22:30 UT E: 3/15-3/20: full data coverage ESA: A: 3/15-3/20: full data coverage B: 3/15-3/20: data gaps 3/15 12:00 - 3/15 14:25 UT, 3/15 16:20 - 3/17 04:40 UT C: 3/15-3/20: data gaps 3/15 15:50 - 3/17 00:02 UT D: 3/15-3/20: data gaps 3/17 17:12 - 3/17 23:00 UT E: 3/15-3/20: full data coverage SST: A: 3/15-3/20: full data coverage B: 3/15-3/20: data gaps 3/15 12:00 - 3/15 14:25 UT, 3/15 16:20 - 3/17 04:40 UT, 3/19 12:20 - 3/19 13:44 UT, no data yet after 03/19 16:25 UT C: 3/15-3/20: data gaps 3/15 15:50 - 3/17 00:02 UT D: 3/15-3/20: data gaps 3/17 17:12 - 3/17 23:00 UT E: 3/15-3/20: full data coverage EFI fields + filter bank: A: 3/15-3/20: full data coverage B: 3/15-3/20: full data coverage C: 3/15-3/20: full data coverage D: 3/15-3/20: data gaps 3/17 17:12 - 3/17 23:00 UT E: 3/15-3/20: full data coverage SCM fields + filter bank: A: 3/15-3/20: full data coverage B: 3/15-3/20: full data coverage C: 3/15-3/20: full data coverage D: 3/15-3/20: data gaps 3/17 17:12 - 3/17 23:00 UT E: 3/15-3/20: full data coverage Triggers and burst data: A: 3/15-3/20: look fine B: 3/15-3/20: look fine C: 3/15-3/20: look fine D: 3/15-3/20: look fine E: 3/15-3/20: look fine ================================= Instrument Ops (Significant events) ------------------------ #--------------------------------------------------------------------------------- # 2009/074 - 2009/075 Sunday, March 15 #--------------------------------------------------------------------------------- BGS_11M THEMIS_B 9.500 2009/074 14:11:00 2009/074 14:51:00 TR9 # Data dump [GL] - no problems after 232 min shadow (06:46 - 10:38) AGO_9M THEMIS_B 9.500 2009/074 15:03:00 2009/074 15:20:00 TR10 # Data dump [GL] USNAU_13M THEMIS_C 9.400 2009/074 15:48:00 2009/074 16:45:00 TR9 # Data dump [GL] - no problems after 162 min shadow (07:46 - 10:28) - station had wrong uplink polarization initially #--------------------------------------------------------------------------------- # 2009/075 - 2009/076 Monday, March 16 Week 12 #--------------------------------------------------------------------------------- BGS_11M THEMIS_D 2.750 2009/076 03:00:35 2009/076 03:20:35 TR2 # SOH check and track - no packets from BAU PCM module BGS_11M THEMIS_D 2.750 2009/076 09:49:25 2009/076 10:09:25 TR2 # SOH check and track - no packets from BAU PCM module #--------------------------------------------------------------------------------- # 2009/076 - 2009/077 Tuesday, March 17 #--------------------------------------------------------------------------------- USNAU_13M THEMIS_D 2.740 2009/076 12:20:00 2009/076 12:40:00 TR2 # SOH check and track - no packets from BAU PCM module USNAU_13M THEMIS_C 9.400 2009/076 15:27:00 2009/076 16:18:00 TR9 # Data dump [DP] - no power problems after 184 min shadow (07:53 - 10:58) - RCS LV2 status indicator discrepancy BGS_11M THEMIS_D 2.770 2009/076 15:34:00 2009/076 15:44:00 TR2 # SOH check and track [DP] - blind acquisition - no packets from BAU PCM module WLP_11M THEMIS_D 2.750 2009/076 19:10:00 2009/076 20:30:00 TR2 # SOH check and track [DP] - recovery from PCM anomaly - turn of instruments and IDPU - # force BAU cold reset without clear - load tables - RF uplink unstable - antenna may be stuck with wrong polarization AGO_9M THEMIS_D 2.570 2009/076 21:58:00 2009/076 23:28:00 TR2 # SOH check and track [DP] - continuation with PCM recovery and instrument turn-on - total instrument downtime 4.0 h #--------------------------------------------------------------------------------- # 2009/077 - 2009/078 Wednesday, March 18 #--------------------------------------------------------------------------------- BGS_11M THEMIS_D 2.750 2009/077 12:43:30 2009/077 13:03:30 TR2 # SOH check and track - no problems after 84 min shadow (07:14 - 08:39) #--------------------------------------------------------------------------------- # 2009/078 - 2009/079 Thursday, March 19 #--------------------------------------------------------------------------------- BGS_11M THEMIS_B 9.500 2009/078 13:28:00 2009/078 13:58:00 TR10 # Data dump [GL] - no problems after 201 min shadow (07:51 - 11:12) AGO_9M THEMIS_B 9.500 2009/078 14:17:00 2009/078 14:37:00 TR10 # Data dump [GL] HBK_10M THEMIS_B 9.500 2009/078 14:55:00 2009/078 15:14:00 TR9 # Data dump [GL] - no socket connections initially from the station's Enertec unit - no data loss BGS_11M THEMIS_C 9.400 2009/078 16:17:15 2009/078 16:36:40 TR10 # Data dump [JM] - no problems after 193 min shadow (08:06 - 11:20) #--------------------------------------------------------------------------------- # 2009/079 - 2009/080 Friday, March 20 #--------------------------------------------------------------------------------- TDRS_6 THEMIS_D 2.750 2009/079 17:56:00 2009/079 18:03:00 TR2 # SOH check and track - TDRS proficiency test - no problems after 134 min shadow (07:12 - 09:27) DSS_27 THEMIS_B 2.992 2009/079 23:30:00 2009/080 00:00:00 TR2 # Ranging test - LRO 3-way proficiency test - MILA 9M2 and WHS 18M1 shadow track